# PennyLang: Pioneering LLM-Based Quantum Code Generation with a Novel PennyLane-Centric Dataset

### v1

@@ Abstract
Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning. Their potential in aiding quantum software development remains underexplored, particularly for the PennyLane framework—a leading platform for hybrid quantum-classical computing. To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific code samples of quantum circuits and their contextual descriptions, specifically curated to train/fine-tune LLM-based quantum code assistance. Our key contributions are threefold: (1) the automatic creation and open-source release of a comprehensive PennyLane dataset leveraging quantum computing textbooks, official documentation, and open-source repositories; (2) the development of a systematic methodology for data refinement, annotation, and formatting to optimize LLM training efficiency; and (3) a thorough evaluation, based on a Retrieval-Augmented Generation (RAG) framework, demonstrating the effectiveness of our dataset in streamlining PennyLane code generation and improving quantum development workflows. Compared to existing efforts that predominantly focus on Qiskit, our dataset significantly broadens the spectrum of quantum frameworks covered in AI-driven code assistance. By bridging this gap and providing reproducible dataset-creation methodologies, we aim to advance the field of AI-assisted quantum programming, making quantum computing more accessible to both newcomers and experienced developers.
@@
Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), Code Completion and Assistance, Dataset Curation and Annotation, Evaluation Framework
@@ I Introduction
Quantum computing has emerged as a transformative technology, with the potential to solve complex problems that are infeasible for classical systems [1, 2]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [3]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [4, 5, 6, 7, 8].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks, hindering widespread adoption [9, 10]. PennyLane is a leading open-source python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [11]. However, unlike the Qiskit framework [12], which benefits from dedicated AI-driven code assistants [13], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has witnessed a surge in the capabilities of LLMs, which excel at reasoning, contextual understanding, and automated code generation [14, 15, 16]. These models have been successfully leveraged in classical programming [17, 18], improving developer efficiency and reducing syntactical and logical errors. However, their application in quantum programming remains largely underexplored, particularly for PennyLane. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for its refinement, annotation, formatting, and evaluation to facilitate LLM training.
State-of-the-Art and Limitations: While several studies have explored LLM-driven code generation for classical programming tasks [19, 20], research on quantum-specific datasets remains scarce. The most notable effort in this domain is the Qiskit Code Assistant [13], which leverages a Qiskit-specific dataset for AI-assisted quantum code generation. However, its hardware-centric approach does not extend to PennyLane, which specializes in variational quantum circuits and quantum machine learning applications [21].
@@
Our Approach and Contributions: In this work, we introduce PennyLang to address the gap in LLM-assisted quantum code generation for pennylane through the following novel contributions. An overview of our contributions is presented in Fig. 1.
@@
    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, two quantum computing textbooks, and the official PennyLane documentation. Each sample is meticulously annotated with explanatory comments and contextual descriptions, facilitating more effective LLM training for quantum programming.
@@
    Systematic Data Refinement and Annotation: We develop a robust methodology for dataset cleaning, eliminating non-PennyLane snippets and enriching the samples with structured metadata. To further enhance usability, we employ the GPT-4o API to convert code samples into an instruction-query format, providing clear guidance for model training in a structured, task-oriented manner.
@@
Summary of Key Results: RAG-enhanced quantum code generation significantly improved performance across all models. GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context. These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability.
Paper Organization: Sec. II provides background information and reviews related work. Sec. III outlines our methodology for dataset collection, cleaning, and annotation along with the tokenization and formatting pipeline used for preprocessing. Sec. IV presents experimental results, including an evaluation of our dataset using a RAG-based framework. Sec. V presents some case studies Finally, Sec. VI concludes the paper with a summary of findings and future research directions.
@@ II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training
@@
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [31] and automated code generation [32], their application in quantum software development remains under-explored. Existing studies on AI-driven code assistants primarily focus on classical programming languages such as Python, Java, and C++, or quantum computing frameworks like Qiskit [13]. However, the PennyLane framework [11], widely used for hybrid quantum-classical computing, lacks dedicated research efforts in AI-assisted quantum programming.
@@ II-B LLMs for Quantum Programming
Despite these advancements, challenges persist:
@@
    Limited Training Data: Quantum programming lacks large-scale, high-quality datasets.
@@
    Complexity of Quantum Circuits: Unlike classical code, quantum programs require understanding quantum mechanics.
@@
    Hybrid Workflows: PennyLane integrates classical and quantum processing, requiring LLMs to learn hybrid paradigms.
These challenges underscore the need for PennyLane-specific datasets for LLM training, as none currently exist, reinforcing the importance of our research.
@@
Research Area Existing Works Our Contributions
@@ III-A3 Scraping Code from PennyLane Documentation
TABLE III: Tokenized Input and Attention Masking
@@ III-A4 Dataset Summary and Statistics
TABLE IV: Dataset Composition
@@ III-A6 Implications for AI-Driven PennyLane Code Generation
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework
@@
TABLE VI: LLMs Used in the RAG-Based Evaluation Framework
@@ III-C1 Retrieval-Augmented Generation (RAG) Framework
To enhance LLM-based quantum code generation, we employ a Retrieval-Augmented Generation (RAG) methodology that integrates retrieval-based context with language model inference. This hybrid approach significantly improves:
  • Accuracy: By grounding responses in retrieved dataset samples, reducing model hallucination.
  • Adherence to PennyLane’s best practices: Ensuring generated quantum code aligns with official standards.
  • Response Quality: Structured retrieval enriches contextual understanding, improving coherence and correctness.
@@ III-C2 RAG Implementation using LangChain
LangChain is utilized to construct a modular and efficient RAG pipeline. The framework consists of the following core components:
  1. 1.
    Dataset Loading: PennyLane dataset samples are extracted from curated dataset structured in intruction-response format.
  2. 2.
    Embedding Generation: Each sample is tokenized and converted into high-dimensional vector embeddings using OpenAIEmbeddings.
  3. 3.
    Vector Storage and Retrieval: The embeddings are stored in a Chroma vector database for fast retrieval.
  4. 4.
    Prompting and Response Generation:
    • When a query is submitted, Maximum Marginal Relevance (MMR) search fetches the most relevant samples in RAG-augmented input.
    • Retrieved context is formatted into a structured prompt.
    • The LLM model generates responses using both RAG-augmented and non-RAG prompts.
  5. 5.
    Evaluation: The responses are compared using a second LLM (GPT-4o Mini), which scores them based on:
    • Functionality: Logical correctness of the quantum circuit.
    • Syntax: Validity of the PennyLane code.
    • Modularity: Adherence to best practices in structuring quantum workflows.
@@
Figure 3: Performance of GPT 4o mini, Claude, and Qwen With and Without RAG. Mean scores (±1 SD) over 29 trials for GPT 4o mini, Claude, and Qwen under both conditions (blue: With RAG; orange: Without RAG), with individual scores shown as swarmplot markers.
Figure 3 presents the results of our RAG-based evaluation methodology, demonstrating a significant performance improvement with retrieval-augmented responses. Each model was tested with 30 code samples, and responses were evaluated using GPT-4o, which graded them based on syntax, functionality, and modularity. The results indicate that the RAG-enhanced approach consistently outperforms the base LLM, yielding higher overall scores.
III-C3 LLM Models and Configuration
The models used in this framework differ in size, architecture, and inference capabilities. Table VI provides a comprehensive comparison of key characteristics including parameter count, context window size, and quantum-specific training data percentage.
III-C4 Prompt Design for RAG and Non-RAG Evaluation
To ensure fair evaluation, we construct two different prompts—one for RAG-augmented inference and another for direct model inference without retrieval.
  • Non-RAG Prompt:
    Base LLM Prompt (Without RAG) "You are an advanced assistant for quantum computing tasks using PennyLane. Provide a response to the user’s question. Make sure to follow best practices and optimize for efficiency. If you are unsure, state that clearly."
  • RAG-Enhanced Prompt:
    RAG-Enhanced Prompt with Context Retrieval "You are an advanced assistant for quantum computing tasks using PennyLane. Use the following pieces of retrieved context to answer the question. Use multiple retrieved code snippets whenever necessary to improve accuracy. If a single snippet does not provide enough information, retrieve and combine relevant parts from multiple sources. Make sure to follow best practices and optimize for efficiency. If you don’t know the answer, say that you don’t know." "{context}"
III-C5 Evaluation Criteria and Scoring
To assess the effectiveness of RAG-based retrieval, generated responses are evaluated using GPT-4o with the following scoring prompt:
Evaluation Prompt for GPT-4o You are an AI assistant tasked with evaluating two different answers to the same question. Question: {question} Answer with RAG: {rag_answer} Answer without RAG: {non_rag_answer} Please provide a score from 1 to 30 for each answer based on: - Functionality (correctness of quantum operations) - Syntax (validity of PennyLane code) - Modularity (code organization and efficiency) Justify your scoring.
III-C6 Findings and Impact of RAG Integration
After evaluating multiple quantum programming queries using the RAG vs. Non-RAG framework, we observe: Based on the evaluation results, RAG-enhanced responses consistently outperformed non-RAG responses across all tested models—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5. On average, RAG integration improved response scores by 11.67% in GPT-4o Mini, 7.69% in Claude 3.5, and 14.38% in Qwen 2.5. The improvement was most noticeable in functionality and syntax, as retrieved context helped prevent incorrect quantum operations and improved PennyLane-specific syntax accuracy. Claude 3.5 Sonnet demonstrated higher coherence and structured responses, while GPT-4o Mini produced concise yet effective outputs. Qwen 2.5 benefited the most from RAG, showing a 14.38% increase in performance, likely due to its stronger reliance on retrieved context for accurate quantum code generation. These results highlight the effectiveness of retrieval-augmented methods in enhancing quantum code generation accuracy and adherence to best practices.
@@ IV-A Response Lengths Across Feature Groups
Figure 5: Box plot of response lengths across different feature groups. Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions
@@ VI Conclusion
This paper presents a dedicated PennyLane dataset and a systematic methodology for preparing quantum code snippets for LLM training. Our open-source release aims to empower researchers and developers to build advanced quantum code assistants that reduce onboarding difficulties and accelerate development in hybrid quantum-classical systems. Ongoing work includes expanding the dataset with more domain-specific samples, conducting broader evaluations with different LLMs, and exploring cross-framework synergy for universal quantum code generation.

### v2

@@ Abstract
Large Language Models (LLMs) offer remarkable capabilities in code generation, natural language processing, and domain-specific reasoning. However, their application in quantum software development remains underexplored, particularly for PennyLane—a leading framework for hybrid quantum-classical computing. To address this gap, we introduce a novel, high-quality dataset comprising 3,347 PennyLane-specific quantum code samples and contextual descriptions, specifically curated to support LLM training and fine-tuning for quantum code assistance. Our contributions are threefold: (1) the automatic construction and open-source release of a comprehensive PennyLane dataset derived from textbooks, official documentation, and open-source repositories; (2) a structured methodology for data curation, annotation, and formatting to enhance LLM usability and relevance; and (3) a rigorous evaluation of code generation capabilities using both baseline Retrieval-Augmented Generation (RAG) and a GraphRAG-enhanced pipeline. Using the PennyLang framework, we demonstrate that GraphRAG, when applied to a GPT-4o Mini model, substantially outperforms standard prompting and baseline RAG. Accuracy improves from 20.5% (without RAG) to 58.2% with GraphRAG, showcasing its effectiveness in reducing hallucinations and improving code correctness in quantum programming tasks. Compared to prior efforts focused largely on Qiskit, our work expands LLM-based assistance to the PennyLane ecosystem, contributing practical tools and reproducible methodologies for advancing AI-assisted quantum software development.
@@
Quantum Computing, Large Language Models (LLMs), PennyLane Framework, AI-Assisted Quantum Programming, Retrieval-Augmented Generation (RAG), GraphRAG, Code Completion and Assistance, Dataset Curation, Evaluation Framework
@@ I Introduction
Quantum computing has emerged as a transformative technology with the potential to solve complex problems that are infeasible for classical systems [1, 2, 3]. Breakthroughs such as Google’s demonstration of “quantum supremacy” [4]—where a quantum processor performed a calculation in seconds that would take thousands of years on a classical supercomputer—highlight its disruptive capabilities in domains ranging from cryptography and quantum chemistry to machine learning and optimization [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
Research Challenges: Despite these advances, programming quantum systems is challenging due to the specialized nature of existing frameworks and benchmarks, hindering widespread adoption [16, 17, 18]. PennyLane is a leading open-source python framework designed for hybrid quantum-classical computing, allowing seamless integration of quantum circuits with machine learning workflows [19]. However, unlike the Qiskit framework [20, 21], which benefits from dedicated AI-driven code assistants [22], PennyLane lacks equivalent tools to assist developers—particularly newcomers—in writing optimized quantum code.
In parallel, the field of Natural Language Processing (NLP) has seen rapid advancements in Large Language Models (LLMs), which now excel at reasoning, contextual understanding, and automated code generation [23, 24, 25]. These models have shown success in classical programming [26, 27], improving developer productivity and reducing syntax and logic errors. However, their application in quantum programming—especially within the PennyLane ecosystem—remains underexplored. This motivates our work: curating a high-quality, PennyLane-centric dataset and establishing a structured methodology for refinement, annotation, and evaluation, ultimately enabling LLM-based quantum code generation and debugging.
State-of-the-Art and Limitations: While LLMs have been benchmarked extensively for classical programming [28, 29], their performance in quantum domains is less understood. Qiskit HumanEval and Qiskit Code Assistant [22] represent early attempts to extend LLM capabilities into quantum programming, but they focus primarily on Qiskit—a hardware-centric framework. PennyLane, in contrast, is designed for variational quantum algorithms and quantum machine learning [30], requiring a different kind of LLM support and training corpus.
@@
Our Approach and Contributions: In this work, we introduce PennyLang, a large-scale, high-quality dataset tailored for PennyLane code generation, along with a rigorous evaluation pipeline. In addition to a baseline RAG framework, we propose and test the GraphRAG framework [32], which leverages structured code relationships to improve the quality and relevance of retrieval during generation. Our contributions are summarized in Fig. 1 and detailed below:
@@
    Comprehensive PennyLane Dataset: We curate a dataset of 3,347 PennyLane-specific code samples sourced from GitHub repositories, quantum computing textbooks, and official PennyLane documentation. Each sample is enriched with contextual descriptions and annotations to support instruction tuning.
@@
    Systematic Data Refinement and Annotation: We present a robust methodology for dataset cleaning and metadata enrichment. Using the GPT-4o API, code snippets are transformed into instruction-query pairs, improving compatibility with LLM training pipelines.
@@
  4. 4.
    GraphRAG-Enhanced Evaluation Framework: We extend the baseline RAG framework with GraphRAG, which introduces a graph-based context retrieval mechanism for code augmentation. GraphRAG models contextual and structural relationships between quantum code samples, enabling more accurate and consistent retrievals. We evaluate this setup using GPT-4o Mini model, comparing GraphRAG, and no-RAG scenarios in a controlled setting.
Summary of Key Results: RAG-enhanced quantum code generation improves performance across all models. GPT-4o Mini showed an 11.67% gain in accuracy and modularity, Claude 3.5 Sonnet improved by 7.6%, enhancing coherence, while Qwen 2.5 saw the highest boost at 14.38%, benefiting from retrieved context. These results confirm RAG’s effectiveness in optimizing PennyLane quantum code generation by improving functionality, syntax, and adaptability. Our results from GraphRAG significantly outperforms both baseline RAG and vanilla prompting. When applied to GPT-4o Mini, GraphRAG boosts accuracy from 20.5% (no RAG) to 58.2%, marking a 37.7% absolute improvement in correct solutions across PennyLang dataset, confirming the benefit of structured retrieval in quantum programming tasks.
Paper Organization: Sec. II reviews related work and background on quantum programming and RAG. Sec. III outlines our dataset construction, annotation methodology, and the GraphRAG retrieval pipeline. Sec. IV presents experimental results comparing model performance across prompting settings. Sec. V discusses case studies and qualitative insights. Sec. VI concludes with future directions for dataset-driven quantum code assistance.
@@ II Background and Related Work
TABLE I: Summary of Qiskit and PennyLane Datasets for LLM Training.
@@
The intersection of LLMs and quantum computing is an emerging research domain. While LLMs have demonstrated impressive capabilities in NLP [41] and automated code generation [42], their application in quantum software development remains underexplored. Most existing studies focus on classical programming languages or quantum frameworks such as Qiskit [22], with little attention to the PennyLane framework [19], despite its popularity in hybrid quantum-classical workflows.
@@ II-B LLMs for Quantum Programming
However, these works face key limitations:
@@
    Lack of High-Quality Data: Quantum programming lacks scalable, curated training datasets.
@@
    Code Complexity: Quantum circuits require LLMs to reason over non-trivial entanglement and measurement logic.
@@
    Hybrid Workflows: PennyLane introduces mixed classical-quantum code patterns that require deeper contextual understanding.
To bridge this gap, we propose the first PennyLane-centric dataset—PennyLang—paired with an advanced evaluation pipeline that incorporates RAG-based retrieval.
II-C Retrieval-Augmented Generation and GraphRAG
Retrieval-Augmented Generation (RAG) [47] has emerged as a popular method to ground LLM outputs using external context. In quantum programming, RAG helps models compensate for limited training data and outdated knowledge by retrieving relevant examples at inference time.
However, baseline RAG systems suffer from semantic drift, irrelevant context injection, and weak code relevance due to shallow vector-based similarity matching. To address this, we employ GraphRAG[32], a graph-enhanced retrieval framework designed to capture structural and functional relationships between quantum code samples. GraphRAG models PennyLang as a graph of interconnected examples, allowing retrievals that are topologically and semantically aligned with the user query.
Compared to traditional RAG, GraphRAG significantly improves context precision and reduces hallucinations in generated code. This leads to large gains in generation accuracy on the PennyLang test dataset.
@@ II-C Retrieval-Augmented Generation and GraphRAG
Research Area Existing Works Our Contributions.
@@ III-A3 Scraping Code from PennyLane Documentation
TABLE III: Tokenized Input and Attention Masking.
@@ III-A4 Dataset Summary and Statistics
TABLE IV: Dataset Composition.
@@ III-A6 Implications for AI-Driven PennyLane Code Generation
TABLE V: PennyLang Components Used in the RAG-Based Evaluation Framework.
@@
TABLE VI: LLMs Used in the RAG-Based Evaluation Framework.
@@ III-C1 Retrieval-Augmented Generation (RAG) Framework
To assess the impact of context-aware prompting on LLM-based quantum code generation, we adopt a Retrieval-Augmented Generation (RAG) framework. This method enhances generation quality by incorporating relevant external context retrieved from our curated PennyLang dataset. RAG helps address common shortcomings in LLMs, such as hallucinations and outdated or incorrect syntax, by anchoring their output in concrete examples from real-world PennyLane programs. This results in improved syntactic validity, functional correctness, and alignment with PennyLane best practices.
@@ III-C2 RAG Implementation using LangChain
Our RAG pipeline is implemented using the LangChain framework, which provides a modular infrastructure for document loading, vector embedding, retrieval, and prompt formatting. The dataset—formatted in an instruction-response structure—is first embedded using the OpenAI Embeddings module. The resulting vectors are stored in a Chroma database for efficient retrieval.
When a user query is submitted (e.g., a quantum programming task or challenge description), we apply Maximum Marginal Relevance (MMR) search to extract the top relevant examples from the dataset. These examples are concatenated into a structured prompt and fed into the LLM for code generation. To ensure fairness and evaluate RAG’s effectiveness, we generate responses both with and without retrieved context using the same query.
III-C3 RAG Evaluation
To benchmark the performance of RAG-augmented LLMs, we evaluate three state-of-the-art models: GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5-7B-Instruct-Turbo. Each model is tested on 29 quantum code generation tasks sampled from the QHack challenge dataset. For each query, both a baseline (non-RAG) and a RAG-enhanced response is generated. We then use GPT-4o to evaluate the outputs according to a rubric comprising three criteria:
  • Functionality: Does the code produce the correct quantum behavior?
  • Syntax: Is the code valid and executable in PennyLane?
  • Modularity: Is the solution well-organized and reusable?
@@ III-C3 RAG Evaluation
Figure 4: Performance of GPT-4o Mini, Claude 3.5 Sonnet, and Qwen 2.5 with and without RAG. Mean scores (±1 SD) over 29 trials are shown for each model. Blue: RAG; Orange: Non-RAG.
Figure 4 summarizes the outcome of these evaluations. RAG-enhanced prompting consistently improves model performance across all criteria. On average, GPT-4o Mini improves by 11.67%, Claude 3.5 by 7.69%, and Qwen 2.5 by 14.38%. These gains highlight how retrieval provides critical context, especially for complex or multi-step quantum circuits.
III-C4 GraphRAG for Advanced Retrieval
To further improve retrieval precision beyond vector-based similarity, we extend our RAG framework using GraphRAG. While the baseline RAG implementation relies solely on Maximum Marginal Relevance (MMR) vector retrieval, GraphRAG incorporates structural relationships between code samples to guide context selection. In this method, the PennyLang dataset is organized into a graph where nodes represent code samples and edges represent functional or structural similarity—such as shared quantum gates, observables, and template usage patterns.
When a query is received, GraphRAG identifies a subgraph around the most semantically similar samples and performs hybrid retrieval by combining graph neighborhood traversal with traditional vector similarity. This enables the selection of more functionally aligned and topologically relevant code snippets, resulting in higher quality context for LLMs.
We evaluate GraphRAG using the same task set as our baseline experiments but restrict the evaluation to the GPT-4o Mini model due to its superior baseline performance. The model was tested on 146 PennyLang test tasks in both no-RAG and GraphRAG settings. The correctness of the generated outputs was determined by whether the solution passed the associated unit tests.
TABLE VII: Performance of GPT-4o Mini: No-RAG vs. GraphRAG (146 tasks).
Metric No-RAG GraphRAG
Correct Solutions 30 85
Incorrect Solutions 116 61
Total Samples 146 146
Accuracy (%) 20.55% 58.22%
As shown in Table VII, the GraphRAG approach yields a dramatic improvement in accuracy—from 20.55% without retrieval to 58.22% with graph-enhanced retrieval. This threefold increase highlights the value of graph-based structural reasoning in quantum programming tasks, particularly for problems involving complex circuit construction or uncommon PennyLane APIs.
Together, these results establish GraphRAG as a significant enhancement to conventional RAG pipelines and demonstrate its practical utility in improving the performance of LLMs on challenging real-world quantum programming benchmarks.
@@ IV-A Response Lengths Across Feature Groups
Figure 5: Box plot of response lengths across different feature groups. Each category represents a specific quantum computing feature set, with response lengths exhibiting variability and outliers, particularly in core operations, measurement, and mathematical functions.
@@ VI Conclusion
In this paper, we introduce PennyLang, a high-quality, instruction-response dataset tailored for PennyLane-based quantum code generation. We propose a robust data collection and annotation methodology, drawing from GitHub repositories, textbooks, and official documentation to curate 3,347 diverse and well-structured code samples. To assess the practical utility of the dataset, we develop a Retrieval-Augmented Generation (RAG) evaluation framework and benchmark three state-of-the-art LLMs—GPT-4o Mini, Claude 3.5 Sonnet, and Qwen2.5-7B—under both RAG and non-RAG conditions. Our findings demonstrate consistent performance gains with RAG, especially in functionality and syntax correctness. Building on this, we introduce a GraphRAG extension that models structural relationships across code snippets using graph-based retrieval. When applied to GPT-4o Mini, GraphRAG yields a 2.8$\times$ improvement in accuracy compared to the non-RAG baseline, highlighting its potential for deeper semantic retrieval in quantum programming. We release both the PennyLang dataset and the evaluation framework to foster reproducibility and encourage further research in LLM-assisted quantum development. Future directions include training lightweight instruction-tuned models on this dataset, scaling GraphRAG across different models and quantum frameworks, and expanding toward cross-framework generalization between PennyLane, Qiskit, and Cirq.